OpenStudy (yttrium):

Find the surface area of the solid of revolution generated by revolving about the x-axis the arc of y^2 + 4x = 2lny from y = 1 to y = 3.

OpenStudy (yttrium):

answer this please.

OpenStudy (anonymous):

\[\int\limits_{}^{}2*\Pi*y*\sqrt{dy^2+dx^2}\]

OpenStudy (anonymous):

surface area integral

OpenStudy (yttrium):

I'm done with those things, but the part making it difficult for me is the integration part. the sqrt(y^4 + y^2 + 1)

OpenStudy (anonymous):

\[\int\limits_{1}^{3}2*\pi*y*\sqrt{1+(dx/dy)^2}dy\]

OpenStudy (yttrium):

Yes. That's make me overthink. Can you help me?

OpenStudy (anonymous):

@dape this is not a surface area integral.

OpenStudy (anonymous):

can you find dx/dy ? 2ydy +4dx=2dy/y (2y-2/y)dy = -4dx dx/dy = 1/(2y) - y/2 (dx/dy)^2 = 1/(4y^2) - 1/2 + y^2/4 1 + (dx/dy)^2 = 1/(4y^2) + 0.5 + y^2/4 so we need to integrate 2pi*sqrt(0.25 + 0.5y^2 + y^4/4) pi * sqrt(1+y^2 + y^4) as you got @Yttrium

OpenStudy (yttrium):

So what shall I do next? :D

OpenStudy (dape):

@Coolsector that's gonna be one hell of an integral.

OpenStudy (anonymous):

well lol what can i do

OpenStudy (dape):

You are gonna have to use elliptic functions to solve it.

OpenStudy (anonymous):

but @dape the integral that you wrote wont give the surface area

OpenStudy (yttrium):

yeah i agree @Coolsector

OpenStudy (dape):

Something seems off, I'm getting a negative answer.

OpenStudy (anonymous):

maybe we need another approach

OpenStudy (yttrium):

btw, how did you get that kind of high smarscore?

OpenStudy (anonymous):

what is it ? lol

OpenStudy (anonymous):

@dape , i know that it is tempting to write the differential dx or dy but we really need to work with ds (the length) http://tutorial.math.lamar.edu/Classes/CalcII/SurfaceArea.aspx look at surface area formulas there

OpenStudy (anonymous):

the hell!! i wrote : so we need to integrate 2pi*sqrt(0.25 + 0.5y^2 + y^4/4) pi * sqrt(1+y^2 + y^4) but it should be 2pi*sqrt(0.25 + 0.5y^2 + y^4/4) pi * sqrt(1+2y^2 + y^4)

OpenStudy (anonymous):

so pi sqrt((y^2+1)^2) pi(y^2+1) !!!

OpenStudy (anonymous):

that is easy to integrate.. just a calculation mistake

OpenStudy (anonymous):

so the surface area is pi * (y^3/3 + y) from 1 to 3 which is 28 and 2/3 times pi

OpenStudy (anonymous):

 first we need to write the formula, since we have the notation about the x-axis, surface area is: on the board then we pick the 2nd definition of ds since we are given 1 <= y <= 3 as the bounds.